OpenStudy (anonymous):

if sinA=-7/25 where 180

hartnn (hartnn):

what A did u get ?

OpenStudy (anonymous):

-16.26020471 and then i substitute it to tan(A/2) which gave me -1/7 but i doubted my answer because of the 180<A<270, how will i do this? is my answer correct??

hartnn (hartnn):

-16.2 is in 4th quadrant, not in 3rd

OpenStudy (anonymous):

ow.. so there is another value for A? which will be in the 3rd quadrant?

hartnn (hartnn):

yes.

hartnn (hartnn):

which i m trying to find

OpenStudy (anonymous):

did you got it?? how??

hartnn (hartnn):

sin(180-(-16.2))=sin(-16.2)=sin(196.2) so A=196.2

hartnn (hartnn):

got that^ ?

OpenStudy (anonymous):

waa!!.. -7 is the answer?? how did u arrived with the sin(180-(-16.2))?

hartnn (hartnn):

because sin(180-B)=sin B so i had sin(-16.2)=-7/25 i wanted to bring the angle in 3rd quadrant, so i used sin(180-B)=sin B

OpenStudy (anonymous):

but when 270<A<360, is my answer correct?? -1/7?

hartnn (hartnn):

yup. 4th quadrant.

OpenStudy (anonymous):

owh.. thanks again..! :)) i have more questions.. i'll just post it :))